OpenStudy (anonymous):

I can't remember the formula for this question. The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference.

myininaya (myininaya):

@Kit_Kat_Nat_<3 I'm testing something. Is this question in the feedback section or the math section? Just used the slope formula by the way.

OpenStudy (camerondoherty):

Feedback...

Nnesha (nnesha):

this is in feedback section

OpenStudy (anonymous):

Oh. Whoops. I thought I was posting this in the math section, sorry, @myininaya

myininaya (myininaya):

Oh its okay. I'm not sure anymore either where things are.

OpenStudy (anonymous):

Lol. I clicked on a question in here and forgot to switch back to math.

myininaya (myininaya):

\[d=\frac{a_{10}-a_1}{10-1}\] <this is what I meant by slope formula

OpenStudy (anonymous):

I haven't ever used that formula before, do you know what the formula is for an arithmetic sequence is? :)

OpenStudy (compassionate):

I don't know what to believe anymore @myininaya

myininaya (myininaya):

\[a_{10}=a_1+(10-1)d \\ a_{10}-a_{1}=(10-1)d \\ \frac{a_{10}-a_{1}}{10-1}=d\] Should be familiar with this equation ...

Nnesha (nnesha):

yes for